The temperature of a particular country

n

 is the number of the month, starting at n

=


n=0

 for January. The temperature in the country is greater than 40 o

C

oC

 for months Jun, July, August, and September. It is less than 40 o

C

oC

 for months January, February, March, April, November and December, and 40 o

C

oC

 for months May and October. If a

a

 is a positive integer, then which of the following options might be correct

T

(

n

)

=

a

(

n

4

)

(

n

9

)

40

T(n)=a(n−4)(n−9)−40

 T

(

n

)

=

a

(

4

n

)

(

n

9

)

40

T(n)=a(4−n)(n−9)−40

 T

(

n

)

=

a

(

4

n

)

(

9

n

)

+

40

T(n)=a(4−n)(9−n)+40

 T

(

n

)

=

a

(

4

n

)

(

n

9

)

+

40

T(n)=a(4−n)(n−9)+40

 The temperature in December is 40

14

a

40−14a

.

 The temperature in April is 40

3

a

40−3a

.


1
Expert's answer
2021-09-29T14:50:25-0400
"T(0)<40, T(1)<40, T(2)<40, T(3)<40,"

"T(4)=40, T(5)>40, T(6)>40, T(7)>40,"

"T(8)>40, T(9)=40, T(10)<40, T(11)<40."

a. "T(n)=a(n\u22124)(n\u22129)\u221240"

False


"T(4)=a(4\u22124)(4\u22129)\u221240=-40<0, a>0"

b. "T(n)=a(4-n)(n\u22129)\u221240"

False


"T(4)=a(4-4)(4\u22129)\u221240=-40<0, a>0"

c. "T(n)=a(4-n)(9-n)+40"

False


"T(4)=a(4-4)(9-4)+40=40, a>0"

"T(9)=a(4-9)(9-9)+40=40, a>0"

"T(0)=a(4-0)(9-0)+40=36a+40>40, a>0"

d. "T(n)=a(4-n)(n\u22129)+40"


"T(4)=a(4-4)(4-9)+40=40, a>0"

"T(9)=a(4-9)(9-9)+40=40, a>0"

"T(n)<40, a>0,n<4 \ or\ n>9"

"T(n)>40, a>0,4<n<9"

Might be the correct representation


e. The temperature in December is "40\u221214a."

"T(11)=40-14a<40, a>0"

Might be the correct representation


f. The temperature in April is "40\u22123a."

"T(3)=40-3a<40, a>0"

Might be the correct representation